The velocity of glycerin in a 5 cm diameter pipe is 1.6 m/s. Find the velocity in a
3 cm diameter pipe which connects with it, both pipes flowing full.
1
Expert's answer
2020-11-23T10:32:05-0500

By the definition of the Law of Continuity, we have:


"A_1v_1=A_2v_2,""v_2=\dfrac{A_1v_1}{A_2},""v_2=\dfrac{\dfrac{\pi(0.05\ m)^2}{4}\cdot 1.6\ \dfrac{m}{s}}{\dfrac{\pi(0.03\ m)^2}{4}}=4.4\ \dfrac{m}{s}."

Answer:

"v_2=4.4\ \dfrac{m}{s}."